### v1

# Semiclassical limit of resonance states in chaotic scattering

Abstract
Resonance states in quantum chaotic scattering systems have a multifractal structure that depends on their decay rate. We show how classical dynamics describes resonance states of all decay rates in the semiclassical limit. This result for chaotic scattering systems corresponds to the well-established quantum ergodicity for closed chaotic systems. Specifically, we generalize Ulam’s matrix approximation of the Perron-Frobenius operator, giving rise to conditionally invariant measures of various decay rates. There are many matrix approximations leading to the same decay rate and we conjecture a criterion for selecting the one relevant for resonance states. Numerically, we demonstrate that resonance states in the semiclassical limit converge to the selected measure. Example systems are a dielectric cavity, the three-disk scattering system, and open quantum maps.
Introduction.—The structure of eigenstates in closed quantum systems, which in the classical limit are ergodic, is described by the semiclassical eigenfunction hypothesis [1, 2, 3] and the quantum ergodicity theorem [4, 5, 6, 7, 8, 9]. In the semiclassical limit almost all eigenstates are uniform on the energy surface in phase space.
In contrast, in quantum scattering systems with chaotic dynamics in the classical limit [10, 11, 12, 13], the structure of resonance states is much more complex [14, 15, 16], see Fig. 1. A recent factorization conjecture states that resonance states are composed of a universal factor given by a complex Gaussian random wave model and a factor of classical origin giving a multifractal structure depending on the decay rate [17, 18, 19]. Indeed, it was proven by Nonnenmacher and Rubin that in the semiclassical limit chaotic resonance states are described by some conditionally invariant measures [20]. Such measures are invariant under classical dynamics up to a change of their norm due to escape [21].
Which are the conditionally invariant measures corresponding to a quantum scattering system? For resonance states close to one specific decay rate, the so-called natural decay rate, the answer is given by the natural measure [14]. It is the eigenvector corresponding to the leading eigenvalue of the Perron-Frobenius operator, which describes the time evolution of densities in phase space [22]. The natural measure has been used extensively in dielectric cavities to describe lasing modes [23, 24, 25, 26, 27, 28, 29, 30, 31, 32, 18, 33]. For such systems with partial escape one can describe resonance states of a second decay rate, the so-called inverse natural decay rate, by the inverse natural measure [34, 35, 36]. Poles of resonance states at the natural decay rate appear in the complex energy plane close to the upper end of the spectrum, while for the inverse natural decay they appear close to the lower end [36, 18].
Figure 1: (a) Visual comparison between the average over 500 resonance states of a dielectric cavity of limaçon shape near $\mathrm{Re}\,kR_{\text{cav}}=2000$ with decay rate near $\gamma=0.053$ (top) and the proposed conditionally invariant measure based on ray dynamics (bottom). (b) Comparison on boundary phase space.
Much less is known for resonance states of any other decay rate [37]. For systems with full escape, e.g. the three-disk scattering system, the support of resonance states is given by invariant sets of the classical dynamics [15, 38, 39, 37, 40, 41, 19]. Furthermore, for quantum maps with full escape the measure in the opening (and its preimages) and how it depends on the decay rate was derived in Ref. [15]. The structure of resonance states was also related to short periodic orbits [42, 43, 44, 45, 46], zeta functions [47, 48], and finite-time Lyapunov exponents [49]. Another object of interest are Schur vectors determined from resonance states [50, 51] which have been described by classical densities [52].
There are approximate heuristic approaches for finding conditionally invariant measures describing resonance states, which are fundamentally different for systems with full escape [53, 19] and partial escape [36, 18]. Resonance states in the semiclassical limit come close, but do not converge, to these measures [36, 19]. So it remains an open question which are the conditionally invariant measures corresponding to the semiclassical limit of chaotic resonance states of all decay rates.
In this paper we construct the conditionally invariant measure that describes the structure of chaotic resonance states of a given decay rate in the semiclassical limit. To this end we generalize Ulam’s matrix approximation of the Perron-Frobenius operator. There are many matrix approximations and we conjecture a criterion for selecting the one relevant for resonance states, giving the desired conditionally invariant measure. Numerically, we demonstrate that resonance states in the semiclassical limit converge to the selected measure. A visual comparison is shown for a dielectric cavity (partial escape) in Fig. 1. Further example systems are the three-disk scattering system (full escape), and quantum maps with partial and full escape.
Ulam’s method.—The time evolution of densities in the phase space of a dynamical system is governed by the Perron-Frobenius operator [22]. A matrix approximation of the operator goes back to Ulam [54] and has found many applications [55, 56, 57, 58, 59, 60, 61]. In the context of scattering systems it was applied to maps with full escape [62] and dielectric cavities [31]. There it generates the natural conditionally invariant measure with the natural decay rate. We first describe this approach, often called Ulam’s method, which is generalized to arbitrary decay rates below.
One partitions phase space into $n$ disjoint cells $\{A_{1},\dots,A_{n}\}$, typically a grid of equally sized boxes. For simplicity, we first consider a time-discrete map $T$ on phase space. It defines for each cell $A_{i}$ subregions $A_{ji}\subset A_{i}$, which are mapped to cells $A_{j}$ under $T$. In other words, each subregion $A_{ji}$ is defined by,
$A_{ji}=A_{i}\cap T^{-1}(A_{j})\;.$ (1)
One defines a transition matrix (or stochastic matrix),
$P_{ji}^{\mathcal{L}}=\frac{\mu_{\mathcal{L}}(A_{ji})}{\mu_{\mathcal{L}}(A_{i})}\;,$ (2)
where $\mu_{\mathcal{L}}$ is the uniform Lebesgue measure on phase space. We denote it as the Lebesgue transition matrix $P^{\mathcal{L}}$ to distinguish it from more general transition matrices $P$ to be introduced below. Each element is the ratio of the Lebesgue measure of the subregion $A_{ji}$ to the Lebesgue measure of the whole cell $A_{i}$. From the above definitions follows $\sum_{j}P^{\mathcal{L}}_{ji}=1$, as required for a transition matrix. The matrix $P^{\mathcal{L}}$ is Ulam’s matrix approximation of the Perron-Frobenius operator. Numerically, it is determined by the fraction of trajectories uniformly started within cell $A_{i}$ that goes to cell $A_{j}$. We mention that for closed systems the right leading eigenvector, $\sum_{i}P^{\mathcal{L}}_{ji}\,\mu_{i}=\mu_{j}$, gives a coarse-grained invariant measure $\mu_{i}=\mu(A_{i})$ for all cells $A_{i}$ of the partition [54, 22]. Note that we use a notation for the order of indices common in physics, as e.g. in Refs. [62, 31].
In a scattering system with partial escape one approximates the reflectivity by a matrix $R$. Each element $R_{ji}$ is representative for the reflectivity of the transition $A_{i}\rightarrow A_{j}$. In a system with full escape some of the $R_{ji}$ will be zero. Combining a transition matrix $P$ with the reflectivity matrix $R$ by element-wise multiplication leads to a matrix $P_{ji}R_{ji}$ with the eigensystem,
$\sum_{i}P_{ji}R_{ji}\,\mu_{i}=\text{e}^{-\gamma}\,\mu_{j}\;.$ (3)
Here the right leading eigenvector defines a coarse-grained conditionally invariant measure $\mu_{i}=\mu(A_{i})$ for all cells $A_{i}$ of the partition. The leading eigenvalue $\text{e}^{-\gamma}$ gives the decay rate $\gamma$ of this measure. Note that $\mu_{i}\geq 0$ for all $i$ is ensured by the Perron-Frobenius theorem [63] for the leading eigenvector. Subleading solutions have no relevance here. This approach was introduced in Ref. [62] for maps with full escape using the Lebesgue transition matrix $P^{\mathcal{L}}$.
For time-continuous systems one also has to consider the transition time [16], e.g. between reflections with a billiard boundary, while in Eq. (3) the transition time was implicitly set to 1. The transition time is approximated by a matrix $t$, where each element $t_{ji}$ is representative for the transition $A_{i}\rightarrow A_{j}$. In this case, the exponential factor in Eq. (3) has to be placed on the left hand side, yielding the eigensystem,
$\sum_{i}P_{ji}R_{ji}\text{e}^{\gamma t_{ji}}\,\mu_{i}=\mu_{j}\;,$ (4)
with leading eigenvalue $1$. Here the unknown decay rate $\gamma$ has to be adjusted, such that the leading eigenvalue of the matrix $P_{ji}R_{ji}\text{e}^{\gamma t_{ji}}$ (multiplied element-wise) is indeed $1$. This can be done iteratively [64]. This approach (without adjusting $\gamma$) was introduced in Ref. [31] for dielectric cavities using the Lebesgue transition matrix $P^{\mathcal{L}}$.
In the limit of increasingly finer partitions of phase space the number of cells increases, $n\rightarrow\infty$. One desires that in this limit the coarse-grained conditionally invariant measure converges, which is mathematically an open question. Numerically, using the Lebesgue transition matrix $P^{\mathcal{L}}$ in Eqs. (3) or (4) one finds convergence to the natural measure $\mu_{\text{nat}}$ and the natural decay rate $\gamma_{\text{nat}}$, respectively [62, 31]. Note that the natural measure $\mu_{\text{nat}}$ is typically not determined using a matrix approximation. Instead a long-term time evolution of any smooth density with the Perron-Frobenius operator is done, which is numerically implemented by iterating trajectories and their intensities [14, 23, 16].
Transition matrices for arbitrary decay rates.—The Lebesgue transition matrix $P^{\mathcal{L}}$, Eq. (2), is based on the restriction that the measure of the subregions $A_{ji}$ of a cell $A_{i}$ is given by the uniform Lebesgue measure. This assumption limits the possible coarse-grained measures to the natural measure decaying with $\gamma_{\text{nat}}$.
In order to find conditionally invariant measures of other decay rates, $\gamma\neq\gamma_{\text{nat}}$, we now lift this restriction. We allow for transition matrices $P\neq P^{\mathcal{L}}$ without any restriction on the properties of the measure on the subregions. We define more general transition matrices,
$P_{ji}=\frac{\mu(A_{ji})}{\mu(A_{i})}\;,$ (5)
with arbitrary measures $\mu$ replacing the uniform Lebesgue measure $\mu_{\mathcal{L}}$ in Eq. (2). Note that small deviations from uniformity were previously used in proofs on the convergence for $n\rightarrow\infty$ [55, 56].
The above definition, Eq. (5), is equivalent to using all transition matrices $P$ (i.e. matrices with non-negative elements and $\sum_{j}P_{ji}=1$), which are compatible with the allowed transitions between the cells of the partition encoded in $P^{\mathcal{L}}$, i.e.
$P_{ji}\neq 0\quad\text{only if}\quad P^{\mathcal{L}}_{ji}\neq 0\;.$ (6)
In other words, for a given partition we allow for all possible matrix approximations $P$ of the Perron-Frobenius operator.
For each such transition matrix $P$ one finds from the leading eigenvector of Eq. (3) or (4) a coarse-grained conditionally invariant measure $\mu$ and its decay rate $\gamma$. Decay rates occur over a wide range that depends on the maximal and minimal entries of the reflectivity matrix $R$, the transition times $t$, and the allowed transitions. In particular, a given decay rate $\gamma$ from this range is found for infinitely many transition matrices $P$, which give rise to infinitely many different coarse-grained conditionally invariant measures $\mu$.
Selection criterion for transition matrix.—Which of the above infinitely many coarse-grained conditionally invariant measures for a given decay rate $\gamma$ is relevant for describing quantum mechanical resonance states with this decay rate? We conjecture that it is the measure emerging from the transition matrix $P$ that is closest to $P^{\mathcal{L}}$, in the sense of minimizing the Kullback-Leibler divergence from $P$ to $P^{\mathcal{L}}$,
$d(P^{\mathcal{L}}||P)=\frac{1}{n}\sum_{i}\,\left(-\sum_{j}P^{\mathcal{L}}_{ji}\,\ln\frac{P_{ji}}{P^{\mathcal{L}}_{ji}}\right)\;,$ (7)
evaluated for each cell $A_{i}$ and averaged over all cells. Its lowest value zero occurs for the special case $P=P^{\mathcal{L}}$, which is consistent with $\gamma=\gamma_{\text{nat}}$. For any other $\gamma\neq\gamma_{\text{nat}}$ one finds $d(P^{\mathcal{L}}||P)>0$. Note that the Kullback-Leibler divergence depends on the order of its arguments and that the chosen ordering uses the Lebesgue transition matrix $P^{\mathcal{L}}$ as the reference.
We can derive this selection criterion for locally randomized scattering systems in the semiclassical limit using a local random vector model in each cell $A_{i}$ [64]. It was introduced for the special case of the randomized Baker map with escape, perfectly describing resonance states of all decay rates [65]. Note that resonance states of the deterministic Baker map, however, showed small deviations from those of the randomized Baker map. We attribute this to anomalies due to the discontinuity of the Baker map.
Figure 2: Cell $A_{i}$ of the partition and its subregions $A_{ji}\subset A_{i}$, Eq. (1). (a) Uniform density giving Lebesgue transition matrix $P^{\mathcal{L}}$. (b) Non-uniform density giving a general transition matrix $P$.
For a better intuition it helps to relate the elements $P_{ji}$ of a transition matrix to a density distribution within a cell $A_{i}$, see Fig. 2. For the Lebesgue transition matrix $P^{\mathcal{L}}$, Eq. (2), the measure $\mu_{\mathcal{L}}(A_{ji})$ of each subregion $A_{ji}$ is obtained by integrating a uniform density. Instead, for a general transition matrix $P$, Eq. (5), we allow for a non-uniform density within the cell. Such $P$ give rise to conditionally invariant measures with other decay rates. The selection criterion chooses among the non-uniform densities, which gives rise to the desired decay rate, the one closest to a uniform density.
Using the method of Lagrange multipliers we select the transition matrix $P$ closest to $P^{\mathcal{L}}$ under the constraints on $P$ and with a measure $\mu$ having the given decay rate $\gamma$. This leads to a set of nonlinear equations [64],
$P_{ji}$ $=\frac{P^{\mathcal{L}}_{ji}}{1+\left(y_{j}R_{ji}\text{e}^{\gamma t_{ji}}-y_{i}\right)\mu_{i}}$ $\forall i,j$ (8)
$\sum_{j}P_{ji}$ $=1$ $\forall i$ (9)
$\sum_{i}P_{ji}R_{ji}\text{e}^{\gamma t_{ji}}\,\mu_{i}$ $=\mu_{j}$ $\forall j$ (10)
$\sum_{i}\mu_{i}$ $=1\;,$ (11)
with the unknown variables of interest, $P_{ji}$ and $\mu_{i}$, as well as the unknown Lagrange multipliers $y_{i}$. Numerically, these nonlinear equations can be solved iteratively and we provide Python code [64]. The special case, $\gamma=\gamma_{\text{nat}}$, leads to $P=P^{\mathcal{L}}$, $\mu=\mu_{\text{nat}}$, and all $y_{i}=0$. In general for $\gamma\neq\gamma_{\text{nat}}$, one finds the elements $P_{ji}$ to be quite different from $P^{\mathcal{L}}_{ji}$, in some cases by large factors.
Example systems.—We will use four example systems to demonstrate for the selected measure (i) the convergence in the limit $n\rightarrow\infty$ of finer partitions and (ii) the agreement with resonance states in the semiclassical limit. These examples cover the cases of partial and full escape for a 2D billiard and a map in each case:
  1. (a)
    Dielectric cavity (partial escape) with limaçon shape, deformation $\varepsilon=0.6$, where it is practically fully chaotic, radius $R_{\text{cav}}$, refractive index $n_{\text{r}}=3.3$, and reflection law for a TM polarized mode [18].
  2. (b)
    Three-disk scattering system (full escape) with disks of radius $a$ and center to center distance $2.1a$ [19].
  3. (c)
    Standard map at kicking strength $K=10$, where it is practically fully chaotic, with partial reflectivity $0.2$ in the interval $q\in[0.3,0.6]$ [36].
  4. (d)
    Like (c) but with reflectivity 0 (full escape) [36].
For details on these systems see [64]. We provide Python code to compute $P^{\mathcal{L}}$, $P$, and $\mu$ [64]. In all cases the elements of $P^{\mathcal{L}}$ are determined using $10^{4}$ trajectories per cell of the partition (started on a uniform grid) for up to $n=3200^{2}\approx 10^{7}$ cells.
(i) Convergence in the limit $n\rightarrow\infty$.—We numerically demonstrate the convergence of the coarse-grained measure in the limit $n\rightarrow\infty$ of a finer partition. To this end we compare the coarse-grained measures for $n$ and for $n/4$ cells by integrating them on a $50\times 50$ grid and using the Jensen-Shannon divergence [66] (its square root being a metric). Figure 3 shows that with increasing $n$ the Jensen-Shannon divergence converges to zero. Furthermore, the ratio of two consecutive Jensen-Shannon divergences is bounded from above by a value smaller than $1$. If this continues to hold for increasing $n$, the coarse-grained measures are a contractive sequence, therefore a Cauchy sequence and thus converge in the limit $n\rightarrow\infty$. This is demonstrated for various decay rates $\gamma\in[\gamma_{\text{nat}},\gamma_{\text{inv}}]$ (partial escape) and $\gamma\geq\gamma_{\text{nat}}$ (full escape) for the billiard systems and maps.
Figure 3: Convergence of coarse-grained conditionally invariant measures for increasingly fine partitions with $n$ cells. Shown is the decay of the Jensen-Shannon divergence $d_{\text{JS}}$ between measures from partitions $n$ and $n/4$. (a) Dielectric cavity for $\gamma\in\{0.011\,(\gamma_{\text{nat}}),0.030,0.053,0.090,0.122\,(\gamma_{\text{inv}})\}$. (b) Three-disk scattering system for $\gamma\in\{0.436\,(\gamma_{\text{nat}}),0.6,1.0,1.4,1.8\}$. (c) Standard map with partial escape for $\gamma\in\{0.22\,(\gamma_{\text{nat}}),0.35,0.55,0.75,0.88\,(\gamma_{\text{inv}})\}$. (d) Standard map with full escape for $\gamma\in\{0.25\,(\gamma_{\text{nat}}),0.35,0.5,0.75,1.0\}$. Symbols $\circ$, $+$, $\square$, $\times$, and $\diamond$ are used for increasing $\gamma$.
Figure 4: Convergence of averaged quantum resonance states to proposed conditionally invariant measures with $n=3200^{2}\approx 10^{7}$ for systems and decay rates of Fig. 3. Shown is the decay of the Jensen-Shannon divergence $d_{\text{JS}}$ in the semiclassical limit, i.e. (a, b) increasing wave number $\mathrm{Re}\,k$ or (c, d) matrix size $N$. Details on the used resonance states are given in [64].
(ii) Comparison with resonance states in the semiclassical limit.—We compare Husimi representations of quantum resonance states near some decay rate $\gamma$ to the corresponding proposed measure from the finest partition. In order to be most sensitive, we use the average over 500 resonance states of similar decay rate. The visual comparison is demonstrated in Fig. 1 for the dielectric cavity at one decay rate, for further examples see [64]. Quantitatively, Fig. 4 shows that going further towards the semiclassical limit the Jensen-Shannon divergence (evaluated on a $50\times 50$ grid) converges to zero. This is demonstrated for various decay rates $\gamma\in[\gamma_{\text{nat}},\gamma_{\text{inv}}]$ (partial escape) and $\gamma\geq\gamma_{\text{nat}}$ (full escape) for the billiard systems and maps. This convergence to zero is in contrast to previous approximate approaches for conditionally invariant measures [53, 36], which describe resonance states quite well, but not perfectly [36, 19]. We attribute different absolute values and convergence speeds to properties of the multifractal resonance states. These numerical results for various types of scattering systems give strong support for the conjectured selection criterion for the transition matrix and the corresponding measure.
Remarks.—Let us stress, that for individual resonance states we find just as well a Jensen-Shannon divergence converging to zero [64]. However, one observes a much larger Jensen-Shannon divergence due to their fluctuations. This makes individual resonance states a much less sensitive test for the quality of a conditionally invariant measure [36].
The analysis is presented for right resonance states. It can be straightforwardly extended to left resonance states of billiards or maps with escape.
Instead of conjecturing closeness of the transition matrix $P$ to $P^{\mathcal{L}}$, as proposed here, one could conjecture closeness of $P$ to a transition matrix $P^{\text{nat}}$ given by Eq. (2) with $\mu_{\mathcal{L}}$ replaced by the natural measure $\mu_{\text{nat}}$ [67]. By definition, this works for the natural decay rate. Also for other decay rates we find in the limit $n\rightarrow\infty$ convergence to the same measures as before. An advantage of using $P^{\text{nat}}$ is that we observe faster convergence in the limit $n\rightarrow\infty$. A disadvantage is that one needs an approximation of the natural measure $\mu_{\text{nat}}$ on a much finer scale than the partition.
Outlook.—It is desirable to find a semiclassical derivation for our selection criterion of the transition matrix leading to the proposed measures, as it is possible for locally randomized systems [64].
Extremely long-lived resonance states with $\gamma<\gamma_{\text{nat}}$ exist at finite wavelengths, but not in the semiclassical limit [37, 35, 19]. Correspondingly, we find coarse-grained conditionally invariant measures for $\gamma<\gamma_{\text{nat}}$, however, for finite number $n$ of cells only and not in the limit $n\rightarrow\infty$. This regime of long-lived resonance states will be studied in the future.
Furthermore, it would be interesting to extend the set of examples to systems with scattering in smooth potential, e.g. models of the chaotic ionization of atoms and of resonances in chemical reactions [68, 69]. Also the relation to the recently found properties of Schur eigenstates [52] needs to be investigated, as well as the relation to quantum and classical channels in bipartite many-body systems [70].
Acknowledgements.
We are grateful for discussions with E. Altmann, A. Bäcker, S. Barkhofen, K. Clauß, T. Harayama, M. Richter, P. Schlagheck, A. Shudo, G. Tanner, T. Tél, and T. Weich. Funded by the Deutsche Forschungsgemeinschaft (DFG, German Research Foundation) – 262765445. The authors gratefully acknowledge the computing time on the high-performance computer at the NHR Center of TU Dresden. This center is jointly supported by the Federal Ministry of Education and Research and the state governments participating in the NHR.

### v2

@@
 Resonance states in quantum chaotic scattering systems have a multifractal structure that depends on their decay rate.
-We show how classical dynamics describes resonance states of all decay rates in the semiclassical limit.
+We show how classical dynamics describes this structure for all decay rates in the semiclassical limit.
 This result for chaotic scattering systems corresponds to the well-established quantum ergodicity for closed chaotic systems.
@@
 Example systems are a dielectric cavity, the three-disk scattering system, and open quantum maps.
-Introduction.—The structure of eigenstates in closed quantum systems, which in the classical limit are ergodic, is described by the semiclassical eigenfunction hypothesis [1, 2, 3] and the quantum ergodicity theorem [4, 5, 6, 7, 8, 9].
+Introduction—The structure of eigenstates in closed quantum systems, which in the classical limit are ergodic, is described by the semiclassical eigenfunction hypothesis [1, 2, 3] and the quantum ergodicity theorem [4, 5, 6, 7, 8, 9].
 In the semiclassical limit almost all eigenstates are uniform on the energy surface in phase space.
@@
 Indeed, it was proven by Nonnenmacher and Rubin that in the semiclassical limit chaotic resonance states are described by some conditionally invariant measures [20].
-Such measures are invariant under classical dynamics up to a change of their norm due to escape [21].
+Such measures are invariant under classical dynamics up to a change of their norm due to escape [21, 22].
 Which are the conditionally invariant measures corresponding to a quantum scattering system?
@@
 Poles of resonance states at the natural decay rate appear in the complex energy plane close to the upper end of the spectrum, while for the inverse natural decay they appear close to the lower end [37, 18].
-Figure 1: (a) Visual comparison between the average over 500 resonance states of a dielectric cavity of limaçon shape near $\mathrm{Re}\,kR_{\text{cav}}=2000$ with decay rate near $\gamma=0.053$ (top) and the proposed conditionally invariant measure based on ray dynamics (bottom).
+Figure 1: (a) Visual comparison between the average over 500 resonance states of a dielectric cavity of limaçon shape near $\mathrm{Re}\,kR_{\text{cav}}=5000$ with decay rate near $\gamma=0.053$ (top) and the proposed conditionally invariant measure based on ray dynamics (bottom).
 (b) Comparison on boundary phase space.
@@
 So it remains an open question which are the conditionally invariant measures corresponding to the semiclassical limit of chaotic resonance states of all decay rates.
-In this paper we construct the conditionally invariant measure that describes the structure of chaotic resonance states of a given decay rate in the semiclassical limit.
+In this Letter we construct the conditionally invariant measure that describes the structure of chaotic resonance states of a given decay rate in the semiclassical limit.
 To this end we generalize Ulam’s matrix approximation of the Perron-Frobenius operator.
@@
 Further example systems are the three-disk scattering system (full escape), and quantum maps with partial and full escape.
-Ulam’s method.—The time evolution of densities in the phase space of a dynamical system is governed by the Perron-Frobenius operator [22].
+Ulam’s method—The time evolution of densities in the phase space of a dynamical system is governed by the Perron-Frobenius operator [23].
 A matrix approximation of the operator goes back to Ulam [55] and has found many applications [56, 57, 58, 59, 60, 61, 62].
@@
 There it generates the natural conditionally invariant measure with the natural decay rate.
-We first describe this approach, often called Ulam’s method, which is generalized to arbitrary decay rates below.
+We first describe this approach, often called Ulam’s method, which is later generalized to arbitrary decay rates.
 One partitions phase space into $n$ disjoint cells $\{A_{1},\dots,A_{n}\}$, typically a grid of equally sized boxes.
@@
 It defines for each cell $A_{i}$ subregions $A_{ji}\subset A_{i}$, which are mapped to cells $A_{j}$ under $T$.
-In other words, each subregion $A_{ji}$ is defined by,
+In other words, each subregion $A_{ji}$ is defined by
 $A_{ji}=A_{i}\cap T^{-1}(A_{j})\;.$ (1)
@@
 We mention that for closed systems the right leading eigenvector, $\sum_{i}P^{\mathcal{L}}_{ji}\,\mu_{i}=\mu_{j}$, gives a coarse-grained invariant measure $\mu_{i}=\mu(A_{i})$ for all cells $A_{i}$ of the partition [55, 23].
-Note that we use a notation for the order of indices common in physics, as e.g.
-in Refs.
+Note that we use a notation for the order of indices common in physics, as, e.g., in Refs.
 [63, 32].
@@
 In a system with full escape some of the $R_{ji}$ will be zero.
-Combining a transition matrix $P$ with the reflectivity matrix $R$ by element-wise multiplication leads to a matrix $P_{ji}R_{ji}$ with the eigensystem,
+Combining a transition matrix $P$ with the reflectivity matrix $R$ by elementwise multiplication leads to a matrix $P_{ji}R_{ji}$ with the eigensystem,
 $\sum_{i}P_{ji}R_{ji}\,\mu_{i}=\text{e}^{-\gamma}\,\mu_{j}\;.$ (3)
@@
 Note that $\mu_{i}\geq 0$ for all $i$ is ensured by the Perron-Frobenius theorem [64] for the leading eigenvector.
-Subleading solutions have no relevance here.
-This approach was introduced in Ref.
+Subleading solutions have positive and negative entries.
+Thus they cannot be interpreted as measures and are of no relevance here.
+The above approach was introduced in Ref.
 [63] for maps with full escape using the Lebesgue transition matrix $P^{\mathcal{L}}$.
@@
 In this case, the exponential factor in Eq.
-(3) has to be placed on the left hand side, yielding the eigensystem,
+(3) has to be placed on the left-hand side, yielding the eigensystem,
 $\sum_{i}P_{ji}R_{ji}\text{e}^{\gamma t_{ji}}\,\mu_{i}=\mu_{j}\;,$ (4)
 with leading eigenvalue $1$.
-Here the unknown decay rate $\gamma$ has to be adjusted, such that the leading eigenvalue of the matrix $P_{ji}R_{ji}\text{e}^{\gamma t_{ji}}$ (multiplied element-wise) is indeed $1$.
+Here the unknown decay rate $\gamma$ has to be adjusted, such that the leading eigenvalue of the matrix $P_{ji}R_{ji}\text{e}^{\gamma t_{ji}}$ (multiplied elementwise) is indeed $1$.
 This can be done iteratively [65].
@@
 One desires that in this limit the coarse-grained conditionally invariant measure converges, which is mathematically an open question.
+In general, the chosen partition and functional space influence the convergence, see the recent Ref.
+[62] and references therein.
 Numerically, using the Lebesgue transition matrix $P^{\mathcal{L}}$ in Eqs.
@@
 Instead a long-term time evolution of any smooth density with the Perron-Frobenius operator is done, which is numerically implemented by iterating trajectories and their intensities [14, 24, 16].
-Transition matrices for arbitrary decay rates.—The Lebesgue transition matrix $P^{\mathcal{L}}$, Eq.
+Transition matrices for arbitrary decay rates—The Lebesgue transition matrix $P^{\mathcal{L}}$, Eq.
 (2), is based on the restriction that the measure of the subregions $A_{ji}$ of a cell $A_{i}$ is given by the uniform Lebesgue measure.
@@
 In particular, a given decay rate $\gamma$ from this range is found for infinitely many transition matrices $P$, which give rise to infinitely many different coarse-grained conditionally invariant measures $\mu$.
-Selection criterion for transition matrix.—Which of the above infinitely many coarse-grained conditionally invariant measures for a given decay rate $\gamma$ is relevant for describing quantum mechanical resonance states with this decay rate?
+Selection criterion for transition matrix—Which of the above infinitely many coarse-grained conditionally invariant measures for a given decay rate $\gamma$ is relevant for describing quantum mechanical resonance states with this decay rate?
 We conjecture that it is the measure emerging from the transition matrix $P$ that is closest to $P^{\mathcal{L}}$, in the sense of minimizing the Kullback-Leibler divergence from $P$ to $P^{\mathcal{L}}$,
@@
 Note that the Kullback-Leibler divergence depends on the order of its arguments and that the chosen ordering uses the Lebesgue transition matrix $P^{\mathcal{L}}$ as the reference.
+Figure 2: Cell $A_{i}$ of the partition and its subregions $A_{ji}\subset A_{i}$, Eq.
+(1).
+(a) Uniform density giving Lebesgue transition matrix $P^{\mathcal{L}}$.
+(b) Nonuniform density giving a general transition matrix $P$.
 We can derive this selection criterion for locally randomized scattering systems in the semiclassical limit using a local random vector model in each cell $A_{i}$ [65].
@@
 We attribute this to anomalies due to the discontinuity of the Baker map.
-Figure 2: Cell $A_{i}$ of the partition and its subregions $A_{ji}\subset A_{i}$, Eq.
-(1).
-(a) Uniform density giving Lebesgue transition matrix $P^{\mathcal{L}}$.
-(b) Non-uniform density giving a general transition matrix $P$.
 For a better intuition it helps to relate the elements $P_{ji}$ of a transition matrix to a density distribution within a cell $A_{i}$, see Fig.
@@
 Instead, for a general transition matrix $P$, Eq.
-(5), we allow for a non-uniform density within the cell.
+(5), we allow for a nonuniform density within the cell.
 Such $P$ give rise to conditionally invariant measures with other decay rates.
-The selection criterion chooses among the non-uniform densities, which gives rise to the desired decay rate, the one closest to a uniform density.
+The selection criterion chooses among the nonuniform densities giving the desired decay rate, the one closest to a uniform density.
 Using the method of Lagrange multipliers we select the transition matrix $P$ closest to $P^{\mathcal{L}}$ under the constraints on $P$ and with a measure $\mu$ having the given decay rate $\gamma$.
@@
 In general for $\gamma\neq\gamma_{\text{nat}}$, one finds the elements $P_{ji}$ to be quite different from $P^{\mathcal{L}}_{ji}$, in some cases by large factors.
-Example systems.—We will use four example systems to demonstrate for the selected measure (i) the convergence in the limit $n\rightarrow\infty$ of finer partitions and (ii) the agreement with resonance states in the semiclassical limit.
+Example systems—We will use four example systems to demonstrate for the selected measure (i) the convergence in the limit $n\rightarrow\infty$ of finer partitions and (ii) the agreement with resonance states in the semiclassical limit.
 These examples cover the cases of partial and full escape for a 2D billiard and a map in each case:
@@
 For details on these systems see [65].
-We provide Python code to compute $P^{\mathcal{L}}$, $P$, and $\mu$ [64].
+We provide Python code to compute $P^{\mathcal{L}}$, $P$, and the proposed measure $\mu$ [65].
 In all cases the elements of $P^{\mathcal{L}}$ are determined using $10^{4}$ trajectories per cell of the partition (started on a uniform grid) for up to $n=3200^{2}\approx 10^{7}$ cells.
-(i) Convergence in the limit $n\rightarrow\infty$.—We numerically demonstrate the convergence of the coarse-grained measure in the limit $n\rightarrow\infty$ of a finer partition.
+(i) Convergence in the limit $n\rightarrow\infty$—We numerically demonstrate the convergence of the coarse-grained measure in the limit $n\rightarrow\infty$ of a finer partition.
 To this end we compare the coarse-grained measures for $n$ and for $n/4$ cells by integrating them on a $50\times 50$ grid and using the Jensen-Shannon divergence [67] (its square root being a metric).
@@
 Details on the used resonance states are given in [65].
-(ii) Comparison with resonance states in the semiclassical limit.—We compare Husimi representations of quantum resonance states near some decay rate $\gamma$ to the corresponding proposed measure from the finest partition.
+(ii) Comparison with resonance states in the semiclassical limit—We compare Husimi representations of quantum resonance states near some decay rate $\gamma$ to the corresponding proposed measure from the finest partition.
 In order to be most sensitive, we use the average over 500 resonance states of similar decay rate.
 The visual comparison is demonstrated in Fig.
-1 for the dielectric cavity at one decay rate, for further examples see [64].
+1 for the dielectric cavity at one decay rate.
+Note that resonance states show multifractality on scales larger than a Planck cell only and we have to smooth the measure on this scale for the visual comparison in Fig.
+1(b).
+For further examples see Supplemental Material [65].
 Quantitatively, Fig.
@@
 These numerical results for various types of scattering systems give strong support for the conjectured selection criterion for the transition matrix and the corresponding measure.
-Remarks.—Let us stress, that for individual resonance states we find just as well a Jensen-Shannon divergence converging to zero [64].
+Remarks—Let us stress that for individual resonance states we find just as well a convergence to the proposed measure [65], as expected by the factorization conjecture [17, 18, 19].
 However, one observes a much larger Jensen-Shannon divergence due to their fluctuations.
 This makes individual resonance states a much less sensitive test for the quality of a conditionally invariant measure [37].
-The analysis is presented for right resonance states. It can be straightforwardly extended to left resonance states of billiards or maps with escape.
+The analysis is presented for right resonance states.
+It can be straightforwardly extended to left resonance states of billiards or maps with escape.
+This should allow for predicting the structure of the left-right Husimi representation [44].
 Instead of conjecturing closeness of the transition matrix $P$ to $P^{\mathcal{L}}$, as proposed here, one could conjecture closeness of $P$ to a transition matrix $P^{\text{nat}}$ given by Eq.
@@
 A disadvantage is that one needs an approximation of the natural measure $\mu_{\text{nat}}$ on a much finer scale than the partition.
-Outlook.—It is desirable to find a semiclassical derivation for our selection criterion of the transition matrix leading to the proposed measures, as it is possible for locally randomized systems [64].
+Outlook—It is desirable to find a semiclassical derivation for our selection criterion of the transition matrix leading to the proposed measures, as it is possible for locally randomized systems [65].
+Numerically, it might be of interest to find an alternative method to compute the proposed measures based on iterating trajectories for long times, as it is common for the natural measure.
 Extremely long-lived resonance states with $\gamma<\gamma_{\text{nat}}$ exist at finite wavelengths, but not in the semiclassical limit [38, 36, 19].
@@
 This regime of long-lived resonance states will be studied in the future.
-Furthermore, it would be interesting to extend the set of examples to systems with scattering in smooth potential, e.g.
-models of the chaotic ionization of atoms and of resonances in chemical reactions [68, 69].
+Furthermore, it would be interesting to extend the set of examples to systems with scattering in smooth potential, e.g., models of the chaotic ionization of atoms and of resonances in chemical reactions [69, 70].
 Also the relation to the recently found properties of Schur eigenstates [53] needs to be investigated, as well as the relation to quantum and classical channels in bipartite many-body systems [71].
 Acknowledgements.
-We are grateful for discussions with E.
+Acknowledgments—We are grateful for discussions with E.
 Altmann, A.
@@
 Weich.
-Funded by the Deutsche Forschungsgemeinschaft (DFG, German Research Foundation) – 262765445.
+Funded by the Deutsche Forschungsgemeinschaft (DFG, German Research Foundation)—262765445.
 The authors gratefully acknowledge the computing time on the high-performance computer at the NHR Center of TU Dresden.